# Grounding Foundation Models through Federated Transfer Learning: A General Framework

### 10 Dec 2023 v5

@@ Abstract.
Foundation Models (FMs) such as GPT-4 encoded with vast knowledge and powerful emergent abilities have achieved remarkable success in various natural language processing and computer vision tasks. Grounding FMs by adapting them to domain-specific tasks or augmenting them with domain-specific knowledge enables us to exploit the full potential of FMs. However, grounding FMs faces several challenges, stemming primarily from constrained computing resources, data privacy, model heterogeneity, and model ownership. Federated Transfer Learning (FTL), the combination of federated learning and transfer learning, provides promising solutions to address these challenges. In recent years, the need for grounding FMs leveraging FTL, coined FTL-FM, has arisen strongly in both academia and industry. Motivated by the strong growth in FTL-FM research and the potential impact of FTL-FM on industrial applications, we propose an FTL-FM framework that formulates problems of grounding FMs in the federated learning setting, construct a detailed taxonomy based on the FTL-FM framework to categorize state-of-the-art FTL-FM works, and comprehensively overview FTL-FM works based on the proposed taxonomy. We also establish correspondences between FTL-FM and conventional phases of adapting FM so that FM practitioners can align their research works with FTL-FM. In addition, we overview advanced efficiency-improving and privacy-preserving techniques because efficiency and privacy are critical concerns in FTL-FM. Last, we discuss opportunities and future research directions of FTL-FM.
@@ 1. Introduction
Federated Transfer Learning (FTL) provides promising solutions to address these obstacles for grounding FMs. We refer to the concept of grounding FMs through FTL techniques as FTL-FM. Specifically, FTL-FM provides servers of FMs and clients of DMs with privacy and security defense tools (Li et al., 2023b), federated learning algorithms (Yang et al., 2019; McMahan et al., 2017), and transfer learning approaches (Pan and Yang, 2010) to collaboratively adapt FMs to domain-specific models or augment FMs with domain-specific knowledge. In recent years, a considerable amount of methods across the spectrum of FTL-FM have been proposed to ground FMs. Subsequently, several survey and vision papers were proposed to overview these works, summarize application scenarios, and discuss potential challenges and solutions (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a; Zhuang et al., 2023; Chen et al., 2023a). However, these works either focus on specific techniques of customizing FMs without considering distributed scenarios of knowledge transfer (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a) or provide high-level overview of federated LLM scenarios, challenges, and potential solutions (Zhuang et al., 2023; Chen et al., 2023a) without formulating FTL-FM problems and systematically reviewing relevant FTL-FM works. To fill these gaps, we propose an FTL-FM framework that formulates the problems of grounding FMs in the federated learning setting. Based on the FTL-FM framework, we systematically review state-of-the-art FTL-FM works. The main contributions of this paper include:
@@
The rest of the paper is organized as follows. Section 2 overviews related works. Section 3 defines a general framework of FTL-FM and Section 4 presents a taxonomy to categorize state-of-the-art FTL-FM works. Moving forward, Section 5, Section 6, and Section 7 delve into a comprehensive review of current FTL-FM works. Subsequently, Section 8 and Section 9 examine efficiency-improving and privacy-preserving methods, respectively. Section 10 discusses opportunities and future directions. Finally, Section 11 presents the concluding remarks for this work.
@@ 2. Related Work
In this section, we briefly overview the existing survey and vision papers related to grounding foundation models (FMs). These papers either focused on techniques of compressing and customizing FMs without considering distributed scenarios of augmenting and adapting FMs (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a), or broadly reviewed federated foundation models without formulating FTL-FM problems and comprehensively reviewing relevant FTL-FM works (Zhuang et al., 2023; Chen et al., 2023a).
@@
General-purpose FMs typically lack domain knowledge. Enhancing FMs with industry-level or domain-specific knowledge requires a large amount of domain data that are often dispersed among multiple private entities (e.g., corporations).  Chen et al. (2023a) proposed a concept of federated large language models (LLMs) and presented a framework for training LLMs in a federated manner. Technically, they focused on the federated learning scenarios in which each client owns a local LLM, and all clients collaboratively pre-train, fine-tune, or prompt engineer their LLMs.  Zhuang et al. (2023) provided an overview of the combination of FMs and FL. In addition to discussing the motivation, challenges, and opportunities of federated FMs, they reviewed the motivations, challenges, opportunities, and future directions of leveraging FMs to facilitate FL.
@@ 3.2. Formulations of Machine Learning Tasks
Various knowledge transfer approaches have been proposed to achieve each of the three objectives defined in the FTL-FM framework (see Definition 3.1). We overview and formulate representative approaches. These formulations are summarized in Table 1. To better illustrate these formulations, we denote the FM hosted by the server as $f_{\mathcal{W}^{S}_{\text{FM}}}(\cdot)$ parameterized by $\mathcal{W}^{S}_{\text{FM}}$ and DM owned by client $k$ as $g_{\mathcal{W}^{C_{k}}}(\cdot)$ parameterized by $\mathcal{W}^{C_{k}}$.
Table 1. An overview of the formulation of representative federated knowledge transfer approaches.
Objective Formulation of Specific Federated Transfer Learning Loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};(x^{C_{k}},y^{C_{k}}))$ Knowledge Transfer Approach Representative Work
@@
$\ell_{\text{TA}}^{C_{k}}(f_{\mathcal{W}^{S}_{\text{FM}}}([x^{C_{k}},g_{\mathcal{W}^{C_{k}}}(x^{C_{k}})]),y^{C_{k}})$ federated prompt tuning (Lin et al., 2023; Sun et al., 2023)
$\ell_{\text{TA}}^{C_{k}}(g_{\mathcal{W}^{C_{k}}_{\text{H}}}\circ f_{\mathcal{W}^{S}_{\text{FM}}}\circ g_{\mathcal{W}^{C_{k}}_{\text{B}}}(x^{C_{k}}),y^{C_{k}})$ federated split learning  (Li et al., 2023b; Tian et al., 2022)
@@
Optimize Foundation Model $\mathcal{W}^{S}_{\text{FM}}$ $\ell_{\text{TA}}^{C_{k}}(\phi_{\text{H}}\circ f_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}\circ\phi_{\text{B}}(x^{C_{k}}),y^{C_{k}})$ where $\phi_{\text{H}}$ and $\phi_{\text{B}}$ are adapters selected from $f_{\mathcal{W}^{S}_{\text{FM}}}$, and $f_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ is emulator compressed from $f_{\mathcal{W}^{S}_{\text{FM}}}$ federated offsite tuning (Xiao et al., 2023; Fan et al., 2023; WeBank, [n. d.]) (Chua et al., 2023; Khalid et al., 2023)
$\ell_{\text{TA}}^{C_{k}}(f_{\mathcal{W}^{S}_{\text{FM}}}(x^{S}),z^{C_{k}})\text{ where }z^{C_{k}}=g_{\mathcal{W}^{C_{k}}}(x^{S})$ federated knowledge distillation (Yu et al., 2023b)
Optimize $\mathcal{W}^{S}_{\text{FM}}$ and $\mathcal{W}^{C_{k}}$, $k=1,\dots,K$ $\ell_{\text{TA}}^{C_{k}}(g_{\mathcal{W}^{C_{k}}}(x^{C_{k}}),y^{C_{k}})\quad\quad\quad\quad\quad\quad\quad\quad$ $\quad\quad\quad\quad+\ell_{\text{MKD}}^{C_{k}}(f_{\mathcal{W}_{\text{FM}}^{S}}(x^{S}),g_{\mathcal{W}^{C_{k}}}(x^{S}))$ federated co-tuning (Fan et al., 2023)
$\ell_{\text{TA}}^{C_{k}}(\phi_{\mathcal{W}^{G}_{a}}\oplus g_{\mathcal{W}^{C_{k}}_{\text{FM}}}(x^{C_{k}}),y^{C_{k}})$ where $\phi_{\mathcal{W}^{G}_{a}}$ is a global PEFT module, and $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is the local FM and is frozen during training federated parameter- efficient fine-tuning  (Zhang et al., 2023c; Zhao et al., 2023b)  (Cai et al., 2023b; Zhang et al., 2023b)
@@
In addition to model transfer, the server can adapt the general knowledge of the FM to the clients’ domain models by transferring data-level and representation-level knowledge. Federated prompt tuning (Lin et al., 2023; Sun et al., 2023), federated split learning (Li et al., 2023b; Tian et al., 2022) and federated knowledge distillation (He et al., 2020) are three representative approaches. We formulate their objectives of transferring knowledge from the server’s FM to the client $k$’s DM as follows:
@@
(8) federated split learning: $\min_{\mathcal{W}^{C_{k}}_{\text{H}},\mathcal{W}^{C_{k}}_{\text{B}}}\ell_{\text{TA}}^{C_{k}}(g_{\mathcal{W}^{C_{k}}_{\text{H}}}\circ f_{\mathcal{W}^{S}_{\text{FM}}}\circ g_{\mathcal{W}^{C_{k}}_{\text{B}}}(x^{C_{k}}),y^{C_{k}}),$
@@
where [·, ·] stands for concatenation, $g\circ f$ is the function composition such that the outputs of the first function are fed into the second function; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training client $k$’s DM using local data and the knowledge from the FM; $\ell_{\text{KD}}^{C_{k}}$ is the knowledge distillation loss for distilling knowledge of FM to client $k$’s DM based on representations $z^{S}$ and $z^{C_{k}}$ shared between the server and client $k$.
@@
As demonstrated in studies (Lin et al., 2023; Sun et al., 2023), in federated prompt tuning, each client $k$ trains a local prompt generator $g_{\mathcal{W}^{C_{k}}}$ using its own local data and the predictions made by $f_{\mathcal{W}^{S}_{\text{FM}}}$ following Eq.(7). These predictions serve as responses to prompts generated by $g_{\mathcal{W}^{C_{k}}}$, along with the input $x^{C_{k}}$. As illustrated by  (Tian et al., 2022), in federated split learning, the FM typically consists of a head model $g_{\mathcal{W}^{C_{k}}_{\text{H}}}$, a backbone model $f_{\mathcal{W}^{S}_{\text{FM}}}$, and a bottom model $g_{\mathcal{W}^{C_{k}}_{\text{B}}}$. The head and bottom models are deployed on the client side, whereas the backbone model is deployed on the server side. The server and clients collaborate to train the dispersed FM following Eq.(8). The core idea of federated knowledge distillation, as demonstrated by  (He et al., 2020), involves the training of each client’s model $g_{\mathcal{W}^{C_{k}}}$ utilizing local datasets $\mathcal{D}^{C_{k}}$ and the representations $z^{S}$ generated by the FM $f_{\mathcal{W}^{S}_{\text{FM}}}$ following Eq.(9).
@@
To achieve the objective of setting 2⃝, federated offsite tuning (Xiao et al., 2023; Fan et al., 2023; WeBank, [n. d.]) and federated knowledge distillation (Yu et al., 2023b) are two representative approaches. We formulate their objectives as follows:
@@
(10) federated offsite tuning: $\min_{\phi_{\text{H}},\phi_{\text{B}}}\ell_{\text{TA}}^{C_{k}}(\phi_{\text{H}}\circ f_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}\circ\phi_{\text{B}}(x^{C_{k}}),y^{C_{k}}),$
(11) federated knowledge distillation: $\min_{\mathcal{W}^{S}_{\text{FM}}}\ell_{\text{KD}}^{C_{k}}(f_{\mathcal{W}^{S}_{\text{FM}}}(x^{S}),z^{C_{k}})\text{ where }z^{C_{k}}=g_{\mathcal{W}^{C_{k}}}(x^{S}).$
@@
where $\phi_{\text{H}}$ and $\phi_{\text{B}}$ are adapters selected from $f_{\mathcal{W}^{S}_{\text{FM}}}$, and $f_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ is emulator compressed from $f_{\mathcal{W}^{S}_{\text{FM}}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training adapters $\phi_{\text{H}}$ and $\phi_{\text{B}}$ with the help of emulator $f_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$; $x^{S}$ is the data shared by the server and clients, and it serves as the base for knowledge distillation, $\ell_{\text{KD}}^{C_{k}}$ is the knowledge distillation loss for training FM (e.g., $f_{\mathcal{W}^{S}_{\text{FM}}}$) using representations $z^{C_{k}}$ generated by client $k$’s DM $g_{\mathcal{W}^{C_{k}}}$ based on $x^{S}$.
@@
The federated offsite tuning is an extension of Offsite-Tuning (Xiao et al., 2023), which involve a head adaptor $\phi_{\text{H}}$ and a bottom adaptor $\phi_{\text{B}}$ selected from the FM $f_{\mathcal{W}^{S}_{\text{FM}}}$ as well as an emulator $f_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ compressed from $f_{\mathcal{W}^{S}_{\text{FM}}}$. As illustrated in  (Fan et al., 2023; WeBank, [n. d.]; Chua et al., 2023), the server first sends $\phi_{\text{H}}$, $\phi_{\text{B}}$, and $f_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ to all clients before training. Then, each client trains the adaptors with the help of the emulator using its local data following Eq.(10). Next, all clients send their adaptors back to the server, which in turn aggregates clients’ adaptors and plugs the aggregated adaptors back to the FM. In federated knowledge distillation, the FM $f_{\mathcal{W}^{S}_{\text{FM}}}$ at the server is trained by the representations generated by each client’s DM $g_{\mathcal{W}^{C_{k}}}$ following Eq.(11) and optionally trained by server’s local data if available.
To achieve the objective of setting 3⃝, Fan et al. (2023) proposed a potential solution to achieve this objective by leveraging mutual knowledge distillation.
(12) federated co-tuning: $\min_{\mathcal{W}^{C_{k}},\mathcal{W}^{S}_{\text{FM}}}\ell_{\text{TA}}^{C_{k}}(g_{\mathcal{W}^{C_{k}}}(x^{C_{k}}),y^{C_{k}})+\ell_{\text{MKD}}^{C_{k}}(f_{\mathcal{W}_{\text{FM}}^{S}}(x^{S}),g_{\mathcal{W}^{C_{k}}}(x^{S})).$
where $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training client $k$’s DM $g_{\mathcal{W}^{C_{k}}}$ using local data; $f_{\mathcal{W}^{S}_{\text{FM}}}$ and $g_{\mathcal{W}^{C_{k}}}$ learn from each other by minimizing the mutual knowledge distillation loss $\ell^{C_{k}}_{\text{MKD}}$ at the server based on server’s data $x^{S}$.
@@
The special case of objective 3⃝ (as formulated in Eq.(5)) is that clients collaboratively train their local FMs with Parameter-Efficient Fine-Tuning (PEFT) techniques. We formulate the objective of each client $k$ as follows.

### 29 Dec 2023 v7

@@ Abstract.
Foundation Models (FMs) such as GPT-4 encoded with vast knowledge and powerful emergent abilities have achieved remarkable success in various natural language processing and computer vision tasks. Grounding FMs by adapting them to domain-specific tasks or augmenting them with domain-specific knowledge enables us to exploit the full potential of FMs. However, grounding FMs faces several challenges, stemming primarily from constrained computing resources, data privacy, model heterogeneity, and model ownership. Federated Transfer Learning (FTL), the combination of federated learning and transfer learning, provides promising solutions to address these challenges. Recently, the need for grounding FMs leveraging FTL, coined FTL-FM, has arisen strongly in both academia and industry. Motivated by the strong growth in FTL-FM research and the potential impact of FTL-FM on industrial applications, we propose an FTL-FM framework that formulates problems of grounding FMs in the federated learning setting, construct a detailed taxonomy based on the FTL-FM framework to categorize state-of-the-art FTL-FM works, and comprehensively overview FTL-FM works based on the proposed taxonomy. We also establish correspondence between FTL-FM and conventional phases of adapting FM so that FM practitioners can align their research works with FTL-FM. In addition, we overview advanced efficiency-improving and privacy-preserving techniques because efficiency and privacy are critical concerns in FTL-FM. Last, we discuss opportunities and future research directions of FTL-FM.
@@ 1. Introduction
Federated Transfer Learning (FTL) provides promising solutions to address these obstacles for grounding FMs. We refer to the concept of grounding FMs through FTL techniques as FTL-FM. Specifically, FTL-FM provides servers of FMs and clients of DMs with privacy and security defense tools (Li et al., 2023c), federated learning algorithms (Yang et al., 2019; McMahan et al., 2017), and transfer learning approaches (Pan and Yang, 2010) to collaboratively adapt FMs to domain-specific models or augment FMs with domain-specific knowledge. In recent years, a considerable amount of methods across the spectrum of FTL-FM have been proposed to ground FMs. Subsequently, several survey and vision papers were proposed to overview these works, summarize application scenarios, and discuss potential challenges and solutions (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a; Zhuang et al., 2023; Chen et al., 2023). However, these works either focus on specific techniques of customizing FMs without considering distributed scenarios of knowledge transfer (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a) or provide high-level overview of federated LLM scenarios, challenges, and potential solutions (Zhuang et al., 2023; Chen et al., 2023) without formulating FTL-FM problems and systematically reviewing relevant FTL-FM works. To fill these gaps, we propose an FTL-FM framework that formulates the problems of grounding FMs in the federated learning setting. Based on the FTL-FM framework, we systematically review state-of-the-art FTL-FM works. The main contributions of this paper include:
@@
The rest of the paper is organized as follows. Section 2 overviews related works. Section 3 defines a general framework of FTL-FM and Section 4 presents a taxonomy to categorize state-of-the-art FTL-FM works. Moving forward, Section 5, Section 6, and Section 7 comprehensively review FTL-FM works of three settings, respectively. Subsequently, Section 9 and Section 10 examine efficiency-improving and privacy-preserving methods, respectively. Section 11 discusses opportunities and future directions. Finally, Section 12 presents the concluding remarks for this work. The outline of core sections of this work is pictorially illustrated in Figure 2.
Figure 2. Outline of core sections of this work.
@@ 2. Related Work
In this section, we briefly overview the existing survey and vision papers related to grounding foundation models (FMs). These papers either focused on techniques of compressing and customizing FMs without considering distributed scenarios of augmenting and adapting FMs (Gou et al., 2021; Ling et al., 2023; Zhu et al., 2023a), or broadly reviewed federated foundation models without formulating FTL-FM problems and comprehensively reviewing relevant FTL-FM works (Zhuang et al., 2023; Chen et al., 2023).
@@
General-purpose FMs typically lack domain knowledge. Enhancing FMs with industry-level or domain-specific knowledge requires a large amount of domain data that are often dispersed among multiple private entities (e.g., corporations).  Chen et al. (2023) proposed a concept of federated large language models (LLMs) and presented a framework for training LLMs in a federated manner. Technically, they focused on the federated learning scenarios in which each client owns a local LLM, and all clients collaboratively pre-train, fine-tune, or prompt engineer their LLMs.  Zhuang et al. (2023) provided an overview of the combination of FMs and FL. In addition to discussing the motivation, challenges, and opportunities of federated FMs, they reviewed the motivations, challenges, opportunities, and future directions of leveraging FMs to facilitate FL.
@@ 3.2. Formulations of Machine Learning Tasks
Various knowledge transfer approaches have been proposed to achieve each of the three objectives defined in the FTL-FM framework (see Definition 3.1). We overview and formulate representative approaches. These formulations are summarized in Table 1. To better illustrate these formulations, we denote the FM hosted by the server as $f_{\mathcal{W}^{S}_{\text{FM}}}(\cdot)$ parameterized by $\mathcal{W}^{S}_{\text{FM}}$, the DM owned by client $k$ as $g_{\mathcal{W}^{C_{k}}}(\cdot)$ parameterized by $\mathcal{W}^{C_{k}}$, the data owned by the server as $\mathcal{D}^{S}=\{x^{S},y^{S}\}$, and the data owned by client $k$ as $\mathcal{D}^{k}=\{x^{C_{k}},y^{C_{k}}\}$.
Table 1. An overview of formulations of representative federated knowledge transfer approaches.
Objective Formulation of Specific Federated Transfer Learning Loss $\ell^{C_{k}}(\mathcal{W}^{S}_{\text{FM}},\mathcal{W}^{C_{k}};\{x^{C_{k}},y^{C_{k}}\})$ Knowledge Transfer Approach Representative Work
@@
$\ell_{\text{TA}}^{C_{k}}(f_{\mathcal{W}^{S}_{\text{FM}}}([x^{C_{k}},g_{\mathcal{W}^{C_{k}}}(x^{C_{k}})]),y^{C_{k}})$ federated prompt tuning (Sun et al., 2022b; Diao et al., 2023; Li et al., 2023b) (Lin et al., 2023; Sun et al., 2023)
$\ell_{\text{TA}}^{C_{k}}(h_{\mathcal{W}^{\diamond}_{\text{HE}}}\circ f_{\mathcal{W}^{S}_{\text{BA}}}\circ g_{\mathcal{W}^{C_{k}}_{\text{BO}}}(x^{C_{k}}),y^{C_{k}})$ where $h_{\mathcal{W}^{\diamond}_{\text{HE}}},\diamond\in\{S,C_{k}\}$, $f_{\mathcal{W}^{S}_{\text{BA}}}$, and $g_{\mathcal{W}^{C_{k}}_{\text{BO}}}$ are the head, backbone, and bottom models split from the FM $f_{\mathcal{W}^{S}_{\text{FM}}}$ federated split learning  (Zhou et al., 2023; Shen et al., 2023; Li et al., 2023c)  (Xu et al., 2023c; Tian et al., 2022)
@@
Optimize Foundation Model $\mathcal{W}^{S}_{\text{FM}}$ $\ell_{\text{TA}}^{C_{k}}(\phi_{\mathcal{W}^{S}_{a}}\circ\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}\circ\psi_{\mathcal{W}^{S}_{b}}(x^{C_{k}}),y^{C_{k}})$ where $\phi_{\mathcal{W}^{S}_{a}}$ and $\psi_{\mathcal{W}^{S}_{b}}$ are adapters selected from $f_{\mathcal{W}^{S}_{\text{FM}}}$, and $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ is emulator compressed from $f_{\mathcal{W}^{S}_{\text{FM}}}$ federated offsite tuning (Xiao et al., 2023; Fan et al., 2023; WeBank, [n. d.]) (Chua et al., 2023; Khalid et al., 2023)
$\ell_{\text{KD}}^{S}(f_{\mathcal{W}^{S}_{\text{FM}}}(x^{S}),z^{C_{k}})\text{ where }z^{C_{k}}=g_{\mathcal{W}^{C_{k}}}(x^{S})$ federated knowledge distillation (Yu et al., 2023b)
Optimize $\mathcal{W}^{S}_{\text{FM}}$ and $\mathcal{W}^{C_{k}}$, $k=1,\dots,K$ $\ell_{\text{TA}}^{C_{k}}(g_{\mathcal{W}^{C_{k}}}(x^{C_{k}}),y^{C_{k}})\quad\quad\quad\quad\quad\quad\quad\quad$ $\quad\quad\quad\quad+\ell_{\text{FT}}(f_{\mathcal{W}_{\text{FM}}^{S}}(x^{a}),y^{a})+\ell_{\text{FT}}(g_{\mathcal{W}^{C_{k}}}(x^{b}),y^{b})$ where $\{x^{a},y^{a}\}$ and $\{x^{b},y^{b}\}$ are fine-tuning data and their specific forms are algorithm-dependent. federated co-optimizing (Fan et al., 2023; Deng et al., 2023)
$\ell_{\text{TA}}^{C_{k}}(\phi_{\mathcal{W}^{G}_{a}}\oplus g_{\mathcal{W}^{C_{k}}_{\text{FM}}}(x^{C_{k}}),y^{C_{k}})$ where $\phi_{\mathcal{W}^{G}_{a}}$ is a global PEFT module, and $g_{\mathcal{W}^{C_{k}}_{\text{FM}}}$ is the local FM and is frozen during training federated parameter- efficient fine-tuning  (Zhang et al., 2023d; Zhao et al., 2023b)  (Cai et al., 2023b; Zhang et al., 2023c)
@@
In addition to model transfer, the server can adapt the general knowledge of the FM to the clients’ domain models by transferring data-level and representation-level knowledge. Federated prompt tuning (Lin et al., 2023; Sun et al., 2023), federated split learning (Li et al., 2023c; Tian et al., 2022) and federated knowledge distillation (He et al., 2020) are three representative approaches. We formulate their objectives of transferring knowledge from the server’s FM to the client $k$’s DM as follows:
@@
(8) federated split learning: $\min_{\mathcal{W}^{C_{k}}_{\text{HE}},\mathcal{W}^{C_{k}}_{\text{BO}}}\ell_{\text{TA}}^{C_{k}}(h_{\mathcal{W}^{\diamond}_{\text{HE}}}\circ f_{\mathcal{W}^{S}_{\text{BA}}}\circ g_{\mathcal{W}^{C_{k}}_{\text{BO}}}(x^{C_{k}}),y^{C_{k}}),$
[…]
@@
where $h_{\mathcal{W}^{\diamond}_{\text{HE}}}$, $f_{\mathcal{W}^{S}_{\text{BA}}}$, and $g_{\mathcal{W}^{C_{k}}_{\text{BO}}}$ are the head, backbone, and bottom models split from the FM $f_{\mathcal{W}^{S}_{\text{FM}}}$, respectively; [·, ·] stands for concatenation; $f\circ g$ is the function composition such that the outputs of the function $g$ are fed into the function $f$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training client $k$’s DM using local data and the knowledge from the FM; $\ell_{\text{KD}}^{C_{k}}$ is the knowledge distillation loss for distilling knowledge of FM to client $k$’s DM based on representations $z^{C_{k}}$ that are shared between the server and client $k$.
@@
As demonstrated in studies (Sun et al., 2022b; Diao et al., 2023; Li et al., 2023b; Lin et al., 2023; Sun et al., 2023), in federated prompt tuning, each client $k$ trains a local prompt generator $g_{\mathcal{W}^{C_{k}}}$ based on its own local data and the responses made by $f_{\mathcal{W}^{S}_{\text{FM}}}$ following Eq.(7). As illustrated by  (Zhou et al., 2023; Shen et al., 2023; Li et al., 2023c; Xu et al., 2023c; Tian et al., 2022), in federated split learning, the FM $f_{\mathcal{W}^{S}_{\text{FM}}}$ typically is split into a head model $h_{\mathcal{W}^{\diamond}_{\text{HE}}},\diamond\in\{S,C_{k}\}$, a backbone model $f_{\mathcal{W}^{S}_{\text{BA}}}$, and a bottom model $g_{\mathcal{W}^{C_{k}}_{\text{BO}}}$. The backbone model is deployed on the server side, the bottom model is deployed on the client side, while the head model can be deployed on either side. The server and clients collaborate to train the dispersed FM following Eq.(8). The core idea of federated knowledge distillation, as illustrated by  (He et al., 2020), involves the training of each client’s model $g_{\mathcal{W}^{C_{k}}}$ based on local datasets $\mathcal{D}^{C_{k}}$ and the representations $z^{S}$ generated by the FM $f_{\mathcal{W}^{S}_{\text{FM}}}$ following Eq.(9).
@@
To achieve the objective of setting 2⃝, federated offsite tuning (Xiao et al., 2023; Fan et al., 2023; WeBank, [n. d.]; Chua et al., 2023; Khalid et al., 2023) and federated knowledge distillation (Yu et al., 2023b) are two representative approaches. We formulate their objectives as follows:
@@
(10) federated offsite tuning: $\min_{\mathcal{W}^{S}_{a},\mathcal{W}^{S}_{b}}\ell_{\text{TA}}^{C_{k}}(\phi_{\mathcal{W}^{S}_{a}}\circ\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}\circ\psi_{\mathcal{W}^{S}_{b}}(x^{C_{k}}),y^{C_{k}}),$
(11) federated knowledge distillation: $\min_{\mathcal{W}^{S}_{\text{FM}}}\ell_{\text{KD}}^{S}(f_{\mathcal{W}^{S}_{\text{FM}}}(x^{S}),z^{C_{k}})\text{ where }z^{C_{k}}=g_{\mathcal{W}^{C_{k}}}(x^{S}).$
@@
where $\phi_{\mathcal{W}^{S}_{a}}$ and $\psi_{\mathcal{W}^{S}_{b}}$ are the two adapters selected from $f_{\mathcal{W}^{S}_{\text{FM}}}$, and $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ is emulator compressed from $f_{\mathcal{W}^{S}_{\text{FM}}}$; $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training $\phi_{\mathcal{W}^{S}_{a}}$ and $\psi_{\mathcal{W}^{S}_{b}}$ with the help of emulator $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$; $x^{S}$ is the data shared by the server and clients, and it serves as the base for knowledge distillation, $\ell_{\text{KD}}^{S}$ is the knowledge distillation loss for fine-tuning $f_{\mathcal{W}^{S}_{\text{FM}}}$ using representations $z^{C_{k}}$ generated by client $k$’s local DM $g_{\mathcal{W}^{C_{k}}}$ based on $x^{S}$.
@@
In federated offsite tuning (Fan et al., 2023; WeBank, [n. d.]; Chua et al., 2023), the server first sends $\phi_{\mathcal{W}^{S}_{a}}$, $\psi_{\mathcal{W}^{S}_{b}}$, and $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ to all clients before training. Then, each client trains the adaptors $\phi_{\mathcal{W}^{S}_{a}}$ and $\psi_{\mathcal{W}^{S}_{b}}$ with the help of the emulator $\tilde{f}_{\widetilde{\mathcal{W}}^{S}_{\text{FM}}}$ based on its local data following Eq.(10). Next, all clients send their adaptors back to the server, which in turn aggregates clients’ adaptors and plugs the aggregated adaptors back to the FM. In federated knowledge distillation, the FM $f_{\mathcal{W}^{S}_{\text{FM}}}$ at the server is fine-tuned by the representations generated by each client’s DM $g_{\mathcal{W}^{C_{k}}}$ following Eq.(11). In both approaches, $f_{\mathcal{W}^{S}_{\text{FM}}}$ can be optionally fine-tuned by the server’s local data if available.
To achieve the objective of setting 3⃝, the server’s FM and clients’ DMs are co-optimized through sharing knowledge of each other. This co-optimizing objective can be formulated as follows.
[…]
$+\ell_{\text{FT}}(f_{\mathcal{W}_{\text{FM}}^{S}}(x^{a}),y^{a})+\ell_{\text{FT}}(g_{\mathcal{W}^{C_{k}}}(x^{b}),y^{b}).$
where $\ell_{\text{TA}}^{C_{k}}$ is the task loss for training client $k$’s DM $g_{\mathcal{W}^{C_{k}}}$ based on local data; $\ell_{\text{FT}}$ is the loss for fine-tuning $f_{\mathcal{W}^{S}_{\text{FM}}}$ and $g_{\mathcal{W}^{C_{k}}}$ based on knoweldge in the form of $\{x^{a},y^{a}\}$ and $\{x^{b},y^{b}\}$, respectively. The specific forms of $\{x^{a},y^{a}\}$ and $\{x^{b},y^{b}\}$ are algorithm-dependent.
The federated co-optimizing can be implemented in various forms, Fan et al. (2023) proposed FedCoLLM (Fan et al., 2023), a federated mutual knowledge distillation approach that enables $f_{\mathcal{W}^{S}_{\text{FM}}}$ and $g_{\mathcal{W}^{C_{k}}}$ to learn from each other through knowledge distillation. In FedCoLLM, $x^{a}=x^{b}=x^{S}$, $y^{a}=g_{W^{C_{k}}}(x^{S})$, and $y^{b}=f_{\mathcal{W}^{S}_{\text{FM}}}(x^{S})$ where $x^{S}$ is public data. Deng et al. (2023) proposed CrossLM that accomplishes the co-optimization based on a synthetic dataset generated by the server’s FM $f_{\mathcal{W}^{S}_{\text{FM}}}$. In CrossLM, $y^{a}=y^{b}=y^{S}$ and $x^{a}=x^{b}=f_{\mathcal{W}^{S}_{\text{FM}}}(\mathcal{F}(y^{S}))$ where $y^{S}$ is the given labels and $\mathcal{F}(y^{S})$ is the corresponding label-descriptive prompts.
@@
A special case of objective 3⃝, as formulated in Eq.(5), is that clients’ DMs are initialized with FMs and they are trained with Parameter-Efficient Fine-Tuning (PEFT) techniques. We formulate the objective of each client $k$ as follows.